# Does Generative Retrieval Overcome the Limitations of Dense Retrieval?

### v1

@@ 1 Introduction
Advances in deep learning and representation learning (Vaswani et al., 2017; Lee & Toutanova, 2018) have established neural information retrieval (IR) as the dominant paradigm (Mitra et al., 2018; Fan et al., 2022). Within this paradigm, dense retrieval (DR) encodes queries and documents into vectors and measures their similarity through bilinear interactions, enabling efficient vectorized recall and delivering state-of-the-art performance across diverse retrieval tasks (Karpukhin et al., 2020; Khattab & Zaharia, 2020). Recently, driven by generative large language models (LLMs) (Radford et al., 2018; Yang et al., 2025b; Lewis et al., 2020), generative retrieval (GR) has emerged as a new branch of neural IR (Tay et al., 2022; Bevilacqua et al., 2022; Zhuang et al., 2022; Wang et al., 2022; Li et al., 2024; Zeng et al., 2024b). GR directly generates identifiers of relevant documents (docids) for a given query, with corpus knowledge embedded in the model parameters. It typically adopts a sequence-to-sequence architecture trained with cross-entropy loss, while inference relies on constrained decoding to ensure valid docids.
To better understand GR, recent studies have examined its connection to DR. Some research interprets GR as implicitly performing dot-product scoring within an LLM’s parameters and propose a unified framework for similarity computation across both paradigms (Nguyen & Yates, 2023; Wu et al., 2024). Despite this formal unification, the substantial differences in model architecture should not be overlooked: DR is encoder-only, whereas GR employs an autoregressive model with a decoder. This naturally raises the question:
@@
To validate our theoretical analysis empirically, we evaluate standard DR, multi-vector DR (MVDR) (Khattab & Zaharia, 2020; Formal et al., 2021; Li et al., 2023a) and two GR variants following the DSI (Tay et al., 2022) framework on the Natural Questions (NQ) (Kwiatkowski et al., 2019) and MS MARCO (Bajaj et al., 2016) datasets. Under controlled settings, we conduct three studies: (i) By varying DR’s negative sampling and embedding dimension, we evaluate their effects on calibration error and ranking metrics; experimental results show optimization limits due to local normalization and representation limits due to the embedding dimension. (ii) By scaling GR and DR with matched model sizes and training corpus sizes, we observe larger gains for GR, providing preliminary evidence that GR has the potential to overcome DR’s bottlenecks when scaled. (iii) Using a larger model with 14B parameters, we conduct zero-shot and test-time scaling experiments for GR and observe promising performance, further supporting the scaling advantages that GR may obtain.
Overall, our theoretical and empirical results highlight key modeling differences between GR and DR, showing that GR avoids DR’s bottlenecks and has greater potential as an IR paradigm at larger data and model scales. However, our experiments are limited to in-distribution queries, and neither the model nor the data scale is arbitrarily large. In practice, GR does not consistently outperform DR, as its effectiveness depends on factors such as docid design (Bevilacqua et al., 2022; Li et al., 2023b), training data construction (Zhuang et al., 2022), and decoding strategies (Zeng et al., 2024a; Lee et al., 2022). We conclude by discussing these limitations and outlining future directions to close the gap between GR’s theoretical promise and practical performance.
@@ 2 Preliminaries
Problem statement. Let $\mathcal{Q}$ be a set of queries and $\mathcal{D}={d_{1},\ldots,d_{N}}$ a document collection. Let $P^{\star}(d\mid q)$ denote the unknown ground-truth conditional distribution of documents given query $q$. Training pairs $(q,d^{+})$ are drawn from a data distribution $\mathcal{D}_{\text{train}}$, where $d^{+}$ is a relevant document under $P^{\star}(\cdot\mid q)$. The goal of IR is to approximate $P^{\star}(d\mid q)$ using a parametric model $P_{\Theta}(d\mid q)$, ensuring both probabilistic calibration and high ranking quality (Chowdhury, 2010).
Dense retrieval. Let $e_{q}\in\mathbb{R}^{r}$ and $e_{d}\in\mathbb{R}^{r}$ denote the query and document embeddings from encoders $f_{q}$ and $f_{d}$, respectively (Karpukhin et al., 2020; Xiong et al., 2020). The DR score for a pair is computed as their inner product $S(q,d)=e_{q}^{\top}e_{d}$, and the locally normalized (e.g., in-batch) softmax loss is defined accordingly:
@@
Eq. 2 encourages $S(q,d^{+})$ to exceed the scores of negatives within the current candidate pool. In practice, negatives may come from the in-batch sampling (Karpukhin et al., 2020; Khattab & Zaharia, 2020) or hard-negative mining (Xiong et al., 2020; Zhan et al., 2021).
Generative retrieval. Each document has a tokenized docid $y_{1:L}\in\mathcal{V}^{L}$ from a finite vocabulary $\mathcal{V}$ (Tay et al., 2022). The GR training loss is defined by a sequence generation model $p_{\Theta}(y_{t}\mid y_{<t},q)$:
@@ 3.1.1 DR locally normalizes surrogate
The DR objective in Eq. 2 minimizes a surrogate defined on the set $\{d^{+}\}\cup\mathcal{N}(q)$, renormalizing scores via a softmax within $K$ candidates per batch. This makes the learning objective explicitly dependent on the sampled negatives, implying that the negative-sampling scheme (both the size of the candidate set and the quality of the negatives) has a substantial impact on the final performance of DR. Ideally, one would use as negatives the entire set of non-relevant documents, but this is computationally infeasible under realistic resource constraints (Wang & Isola, 2020). This mismatch leads to a calibration gap between the global and local objectives.
@@
Practical mitigations for the calibration gap. Increasing $K$ and mining harder negatives can partially reduce the gap by better approximating the global normalization, and temperature scaling or post-hoc calibration further helps align scores (Xiong et al., 2020; Zhan et al., 2021). Nevertheless, as the corpus size $N$ grows, the $\log(N/K)$ term dominates unless $K$ scales proportionally with $N$, making it increasingly hard for DR to match the true posterior calibration.
@@ 3.2.1 DR exhibits a low-rank bottleneck in relevance representation
DR learns a text-to-embedding mapping and computes relevance through a fixed post-interaction rule, typically a bilinear score such as the inner product $S(q,d)=e_{q}^{\top}e_{d}$. Consequently, all relevance information for a query or a document is compressed into an $r$-dimensional vector (Weller et al., 2025). Formally, DR stacks $m$ query embeddings into $Q\in\mathbb{R}^{m\times r}$ and $N$ document embeddings into $D\in\mathbb{R}^{N\times r}$. The resulting relevance matrix is $S\;=\;QD^{\top}\;\in\;\mathbb{R}^{m\times N}$, which satisfies $\operatorname{rank}(S)\leq r$ regardless of the encoder architecture, as long as the final interaction is bilinear.
By the Eckart-Young-Mirsky theorem (Eckart & Young, 1936; Mirsky, 1960), among all matrices of rank at most $r$, the truncated SVD of any target logit matrix $S^{\star}$ achieves the best Frobenius-norm approximation, with minimal error equal to the sum of squared discarded singular values. We therefore state the following corollary:
@@
Whenever $S^{\star}$ exhibits a heavy spectral tail, a fixed-$r$ DR model inevitably suffers from an irreducible approximation error unless $r$ is increased. Contemporaneous work (Weller et al., 2025) also identifies this limitation of DR, providing detailed proofs and experiments, and argues that late-interaction MVDR models (e.g., ColBERT (Khattab & Zaharia, 2020)) may mitigate the issue. However, we show that MVDR remains subject to a similar upper bound when tokens are grouped into channels (see Appendix D for details).
@@ 3.2.2 GR directly fits the query-document relevance mapping
Theorem 3.4 states that under a fixed bijective docid coding and for in-distribution queries, a sufficiently expressive GR model can approximate the true query–document relevance mapping arbitrarily well (in expected total-variation distance). In other words, with adequate capacity, GR could represents documents, queries, and their relevance relations within the model itself. Note that Theorem 3.4 continues to hold when GR decodes under prefix-constrained decoding (see Appendix C for a detailed proof). Nevertheless, in practice the degree to which GR fits the query-document mapping is affected by several factors, including the quality of the docid tree design and the sufficiency and cleanliness of training data (Tay et al., 2022; Zhuang et al., 2022; Wang et al., 2022). Therefore, Theorem 3.4 is a capacity statement rather than a claim about sample or compute efficiency. It assumes an in-distribution query law and a fixed docid. A highly unbalanced or semantically incoherent docid trie can increase optimization difficulty even under universality, and no guarantee is made for out-of-distribution queries.
@@ 4.1 Experimental setup
We evaluate on two widely used retrieval benchmarks: (i) Natural Questions (NQ)(Kwiatkowski et al., 2019): Real user questions paired with supporting evidence from Wikipedia; and (ii) MS MARCO Passage(Bajaj et al., 2016): Web search queries from Bing with associated relevant passages. We report the calibration metric Brier, computed as the mean squared error between the predicted relevance probability for the top-1 candidate and the ground truth for each query. We also report three retrieval metrics: (i) Hits@$k$, (ii) NDCG@$k$, and (iii) MRR@$k$.
We implement representative systems for DR and GR, deliberately avoiding sophisticated variants to ensure fairness and transparency. For DR, we use: (i) a standard dual encoder with inner-product scoring (referred to as Standard DR), following DPR (Karpukhin et al., 2020); and (ii) a multi-vector late-interaction variant (referred to as MVDR) in the style of ColBERT-v1 (Khattab & Zaharia, 2020). For GR, we adopt two docid designs and follow a DSI-style training/inference pipeline (Tay et al., 2022): (i) codebook docidsconstructed via residual quantization, where each docid is a length-6 sequence of 8-bit code indices (referred to as GR-codebook); and (ii) text docidsthat directly use the document title as the identifier (referred to as GR-text). All GR decoding is prefix-constrained by a trie built from the set of valid docids.
To control for capacity and pretraining, all DR models are built on Qwen3-Embedding-0.6B, and all GR models use Qwen3-0.6B (Yang et al., 2025a). Full details of the experimental setup are provided in Appendix F, and the implementation details for each subsequent experiment are given in Appendix G.
@@ 5 Discussion
Practical challenges of GR. Although GR is theoretically appealing and exhibits demonstrable scaling advantages, it seldom reaches the theoretical optimum in practice, for two main reasons: (i) Noisy or biased supervision (e.g., conflicting relevance labels) and insufficient training can induce an irreducible mismatch between the learned model and the target posterior (Zhuang et al., 2022); and (ii) Prefix-constrained autoregressive decoding is prone to error propagation which means once early tokens deviate, subsequent steps tend to drift (Bevilacqua et al., 2022; Zhang et al., 2024). This issue is exacerbated when the docid design is flawed (e.g., unbalanced hierarchies, suboptimal clustering, or text-based docids that fail to cover document content). Beyond this optimality gap, engineering considerations further limit GR’s practical use: (i) GR’s token-by-token decoding introduces high per-step latency whereas ANN-indexed DR can provide near-instant lookups once the index is built; and (ii) under continual corpus drift, GR often needs retraining or local fine-tuning to accommodate an updated codebook or shifting hierarchical boundaries (Chen et al., 2023; Kishore et al., 2023), whereas DR commonly supports index-only updates.
Potential solutions. We discuss some potential solutions to address the practical challenges of GR. For data noise and undertraining, two complementary directions are promising: (i) treating relevance itself as the pretraining target and pretrain a decoder-only model from scratch on large-scale, noise-controlled $(q,d)$ pairs to directly optimize $\log P(d\mid q)$, similar to some recent works on generative recommendation (e.g., one-rec (Deng et al., 2025)). This is appropriate when relevance is explicitly defined by human rules (e.g., e-commerce query–item (Rajput et al., 2023), ads matching (Fan et al., 2019), FAQ–KB pairs (Sakata et al., 2019)); and (ii) exploiting the world knowledge and reasoning of LLM bases. Specifically, teach the model the semantics and interface of retrieval with light instruction tuning instead of memorizing full-corpus relevance. At inference, execute “retrieval as constrained generation” via constrained decoding. This is suitable when the relevance underlying the retrieval task is already encoded in the pretraining corpus (e.g., Wikipedia or encyclopedic retrieval (Petroni et al., 2020)).
@@ Acknowledgments
This work was funded by the Strategic Priority Research Program of the CAS under Grants No. XDB0680102, the National Natural Science Foundation of China (NSFC) under Grants No. 62472408, 62372431 and 62441229, the National Key Research and Development Program of China under Grants No. 2023YFA1011602, the Dutch Research Council (NWO), under project numbers 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006, and the European Union under grant agreements No. 101070212 (FINDHR) and No. 101201510 (UNITE).
@@ Appendix F Detailed experimental setup
Datasets. We evaluate on two standard retrieval benchmark datasets: (i) Natural Questions (NQ)(Kwiatkowski et al., 2019). This is a collection of real-user questions paired with supporting Wikipedia evidence. We use the official train (313K) and test (7K) splits. To make generative retrieval feasible, we ensure that each test query’s gold document appears in the docid inventory constructed from the training corpus (i.e., the gold docid is seen during training); and (ii) MS MARCO Passage(Bajaj et al., 2016). This is a set of web search queries from Bing with associated passages. We use the passage-ranking subset and sample 1M training pairs and 2K evaluation queries from the official train/test splits. Unlike NQ, we do not enforce the “seen-document” constraint on MS MARCO (because enforcing it would shrink the evaluation set to only few hundred queries).
Models used for comparison. We implement two representative systems for both DR and GR and intentionally avoid complex variants to keep comparisons fair and transparent. For DR, we implement (i) a standard bi-encoder in the spirit of DPR (Karpukhin et al., 2020) with inner-product scoring; and (ii) a multi-vector late-interaction variant like ColBERT v1 (Khattab & Zaharia, 2020). For GR, we implement two varying about the docid design and train/inference follow the DSI-style (Tay et al., 2022): (i) codebook docidsbuilt via residual quantization, each docid is a length-6 sequence of 8-bit code indices; and (ii) textual docidsthat directly use the title as the document identifier. All GR decoding is prefix-constrained by a trie constructed from the set of valid docids.
@@
Training and inference. To control for capacity and pretraining, all DR models are built on Qwen3-Embedding-0.6B, and all GR models use Qwen3-0.6B (Yang et al., 2025a). Unless otherwise noted, we train with the Adam optimizer (Kingma & Ba, 2014) using its default settings. At inference time, DR retrieves top-$k$ candidates using FAISS-based ANN search (Xiong et al., 2020), while GR performs top-$k$ constrained decoding over the docid trie.

### v2

@@ 1 Introduction
Advances in deep learning and representation learning (attention; bert) have established neural information retrieval (IR) as the dominant paradigm (neuralir; pretrainir). Within this paradigm, dense retrieval (DR) encodes queries and documents into vectors and measures their similarity through bilinear interactions, enabling efficient vectorized recall and delivering state-of-the-art performance across diverse retrieval tasks (dpr; colbert). Recently, driven by generative large language models (LLMs) (gpt; qwen2.5; bart), generative retrieval (GR) has emerged as a new branch of neural IR (dsi; seal; dsi-qg; nci; ltgtr; pag). GR directly generates identifiers of relevant documents (docids) for a given query, with corpus knowledge embedded in the model parameters. It typically adopts a sequence-to-sequence architecture trained with cross-entropy loss, while inference relies on constrained decoding to ensure valid docids.
To better understand GR, recent studies have examined its connection to DR. Some research interprets GR as implicitly performing dot-product scoring within an LLM’s parameters and propose a unified framework for similarity computation across both paradigms (gr-dr; gr-mvdr). Despite this formal unification, the substantial differences in model architecture should not be overlooked: DR is encoder-only, whereas GR employs an autoregressive model with a decoder. This naturally raises the question:
@@
To validate our theoretical analysis empirically, we evaluate standard DR, multi-vector DR (MVDR) (colbert; splade; slim) and two GR variants following the DSI (dsi) framework on the Natural Questions (NQ) (nq) and MS MARCO (msmarco) datasets. Under controlled settings, we conduct three studies: (i) By varying DR’s negative sampling and embedding dimension, we evaluate their effects on calibration error and ranking metrics; experimental results show optimization limits due to local normalization and representation limits due to the embedding dimension. (ii) By scaling GR and DR with matched model sizes and training corpus sizes, we observe larger gains for GR, providing preliminary evidence that GR has the potential to overcome DR’s bottlenecks when scaled. (iii) Using a larger model with 14B parameters, we conduct zero-shot and test-time scaling experiments for GR and observe promising performance, further supporting the scaling advantages that GR may obtain.
Overall, our theoretical and empirical results highlight key modeling differences between GR and DR, showing that GR avoids DR’s bottlenecks and has greater potential as an IR paradigm at larger data and model scales. However, our experiments are limited to in-distribution queries, and neither the model nor the data scale is arbitrarily large. In practice, GR does not consistently outperform DR, as its effectiveness depends on factors such as docid design (seal; minder), training data construction (dsi-qg), and decoding strategies (ripor; non-parametric-gr). We conclude by discussing these limitations and outlining future directions to close the gap between GR’s theoretical promise and practical performance.
@@ 2 Preliminaries
Problem statement. Let $\mathcal{Q}$ be a set of queries and $\mathcal{D}={d_{1},\ldots,d_{N}}$ a document collection. Let $P^{\star}(d\mid q)$ denote the unknown ground-truth conditional distribution of documents given query $q$. Training pairs $(q,d^{+})$ are drawn from a data distribution $\mathcal{D}_{\text{train}}$, where $d^{+}$ is a relevant document under $P^{\star}(\cdot\mid q)$. The goal of IR is to approximate $P^{\star}(d\mid q)$ using a parametric model $P_{\Theta}(d\mid q)$, ensuring both probabilistic calibration and high ranking quality (introduction-ir).
Dense retrieval. Let $e_{q}\in\mathbb{R}^{r}$ and $e_{d}\in\mathbb{R}^{r}$ denote the query and document embeddings from encoders $f_{q}$ and $f_{d}$, respectively (dpr; ance). The DR score for a pair is computed as their inner product $S(q,d)=e_{q}^{\top}e_{d}$, and the locally normalized (e.g., in-batch) softmax loss is defined accordingly:
@@
Eq. 2 encourages $S(q,d^{+})$ to exceed the scores of negatives within the current candidate pool. In practice, negatives may come from the in-batch sampling (dpr; colbert) or hard-negative mining (ance; hard-negatives).
Generative retrieval. Each document has a tokenized docid $y_{1:L}\in\mathcal{V}^{L}$ from a finite vocabulary $\mathcal{V}$ (dsi). The GR training loss is defined by a sequence generation model $p_{\Theta}(y_{t}\mid y_{<t},q)$:
@@ 3.1.1 DR locally normalizes surrogate
The DR objective in Eq. 2 minimizes a surrogate defined on the set $\{d^{+}\}\cup\mathcal{N}(q)$, renormalizing scores via a softmax within $K$ candidates per batch. This makes the learning objective explicitly dependent on the sampled negatives, implying that the negative-sampling scheme (both the size of the candidate set and the quality of the negatives) has a substantial impact on the final performance of DR. Ideally, one would use as negatives the entire set of non-relevant documents, but this is computationally infeasible under realistic resource constraints (contrastive-representation). This mismatch leads to a calibration gap between the global and local objectives.
@@
Practical mitigations for the calibration gap. Increasing $K$ and mining harder negatives can partially reduce the gap by better approximating the global normalization, and temperature scaling or post-hoc calibration further helps align scores (ance; hard-negatives). Nevertheless, as the corpus size $N$ grows, the $\log(N/K)$ term dominates unless $K$ scales proportionally with $N$, making it increasingly hard for DR to match the true posterior calibration.
@@ 3.2.1 DR exhibits a low-rank bottleneck in relevance representation
DR learns a text-to-embedding mapping and computes relevance through a fixed post-interaction rule, typically a bilinear score such as the inner product $S(q,d)=e_{q}^{\top}e_{d}$. Consequently, all relevance information for a query or a document is compressed into an $r$-dimensional vector (embedding-rank). Formally, DR stacks $m$ query embeddings into $Q\in\mathbb{R}^{m\times r}$ and $N$ document embeddings into $D\in\mathbb{R}^{N\times r}$. The resulting relevance matrix is $S\;=\;QD^{\top}\;\in\;\mathbb{R}^{m\times N}$, which satisfies $\operatorname{rank}(S)\leq r$ regardless of the encoder architecture, as long as the final interaction is bilinear.
By the Eckart-Young-Mirsky theorem (Eckart–Young–Mirsky1; Eckart–Young–Mirsky2), among all matrices of rank at most $r$, the truncated SVD of any target logit matrix $S^{\star}$ achieves the best Frobenius-norm approximation, with minimal error equal to the sum of squared discarded singular values. We therefore state the following corollary:
@@
Whenever $S^{\star}$ exhibits a heavy spectral tail, a fixed-$r$ DR model inevitably suffers from an irreducible approximation error unless $r$ is increased. Contemporaneous work (embedding-rank) also identifies this limitation of DR, providing detailed proofs and experiments, and argues that late-interaction MVDR models (e.g., ColBERT (colbert)) may mitigate the issue. However, we show that MVDR remains subject to a similar upper bound when tokens are grouped into channels (see Appendix D for details).
@@ 3.2.2 GR directly fits the query-document relevance mapping
Theorem 3.4 states that under a fixed bijective docid coding and for in-distribution queries, a sufficiently expressive GR model can approximate the true query–document relevance mapping arbitrarily well (in expected total-variation distance). In other words, with adequate capacity, GR could represents documents, queries, and their relevance relations within the model itself. Note that Theorem 3.4 continues to hold when GR decodes under prefix-constrained decoding (see Appendix C for a detailed proof). Nevertheless, in practice the degree to which GR fits the query-document mapping is affected by several factors, including the quality of the docid tree design and the sufficiency and cleanliness of training data (dsi; dsi-qg; nci). Therefore, Theorem 3.4 is a capacity statement rather than a claim about sample or compute efficiency. It assumes an in-distribution query law and a fixed docid. A highly unbalanced or semantically incoherent docid trie can increase optimization difficulty even under universality, and no guarantee is made for out-of-distribution queries.
@@ 4.1 Experimental setup
We evaluate on two widely used retrieval benchmarks: (i) Natural Questions (NQ)(nq): Real user questions paired with supporting evidence from Wikipedia; and (ii) MS MARCO Passage(msmarco): Web search queries from Bing with associated relevant passages. We report the calibration metric Brier, computed as the mean squared error between the predicted relevance probability for the top-1 candidate and the ground truth for each query. We also report three retrieval metrics: (i) Hits@$k$, (ii) NDCG@$k$, and (iii) MRR@$k$.
We implement representative systems for DR and GR, deliberately avoiding sophisticated variants to ensure fairness and transparency. For DR, we use: (i) a standard dual encoder with inner-product scoring (referred to as Standard DR), following DPR (dpr); and (ii) a multi-vector late-interaction variant (referred to as MVDR) in the style of ColBERT-v1 (colbert). For GR, we adopt two docid designs and follow a DSI-style training/inference pipeline (dsi): (i) codebook docidsconstructed via residual quantization, where each docid is a length-6 sequence of 8-bit code indices (referred to as GR-codebook); and (ii) text docidsthat directly use the document title as the identifier (referred to as GR-text). All GR decoding is prefix-constrained by a trie built from the set of valid docids.
To control for capacity and pretraining, all DR models are built on Qwen3-Embedding-0.6B, and all GR models use Qwen3-0.6B (qwen3). Full details of the experimental setup are provided in Appendix F, and the implementation details for each subsequent experiment are given in Appendix G.
@@ 5 Discussion
Practical challenges of GR. Although GR is theoretically appealing and exhibits demonstrable scaling advantages, it seldom reaches the theoretical optimum in practice, for two main reasons: (i) Noisy or biased supervision (e.g., conflicting relevance labels) and insufficient training can induce an irreducible mismatch between the learned model and the target posterior (dsi-qg); and (ii) Prefix-constrained autoregressive decoding is prone to error propagation which means once early tokens deviate, subsequent steps tend to drift (seal; term-set-gr). This issue is exacerbated when the docid design is flawed (e.g., unbalanced hierarchies, suboptimal clustering, or text-based docids that fail to cover document content). Beyond this optimality gap, engineering considerations further limit GR’s practical use: (i) GR’s token-by-token decoding introduces high per-step latency whereas ANN-indexed DR can provide near-instant lookups once the index is built; and (ii) under continual corpus drift, GR often needs retraining or local fine-tuning to accommodate an updated codebook or shifting hierarchical boundaries (continual-gr; incdsi), whereas DR commonly supports index-only updates.
Potential solutions. We discuss some potential solutions to address the practical challenges of GR. For data noise and undertraining, two complementary directions are promising: (i) treating relevance itself as the pretraining target and pretrain a decoder-only model from scratch on large-scale, noise-controlled $(q,d)$ pairs to directly optimize $\log P(d\mid q)$, similar to some recent works on generative recommendation (e.g., one-rec (onerec)). This is appropriate when relevance is explicitly defined by human rules (e.g., e-commerce query–item (tiger), ads matching (ads), FAQ–KB pairs (faq)); and (ii) exploiting the world knowledge and reasoning of LLM bases. Specifically, teach the model the semantics and interface of retrieval with light instruction tuning instead of memorizing full-corpus relevance. At inference, execute “retrieval as constrained generation” via constrained decoding. This is suitable when the relevance underlying the retrieval task is already encoded in the pretraining corpus (e.g., Wikipedia or encyclopedic retrieval (kilt)).
@@ Acknowledgments
This work was funded by the National Key Research and Development Program of China under Grants No. 2023YFA1011602, the Strategic Priority Research Program of the CAS under Grants No. XDB0680102, the National Natural Science Foundation of China (NSFC) under Grants No. 62472408, 62372431 and 62441229, the Dutch Research Council (NWO), under project numbers 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006, and the European Union under grant agreements No. 101070212 (FINDHR) and No. 101201510 (UNITE).
@@ Appendix F Detailed experimental setup
Datasets. We evaluate on two standard retrieval benchmark datasets: (i) Natural Questions (NQ)(nq). This is a collection of real-user questions paired with supporting Wikipedia evidence. We use the official train (313K) and test (7K) splits. To make generative retrieval feasible, we ensure that each test query’s gold document appears in the docid inventory constructed from the training corpus (i.e., the gold docid is seen during training); and (ii) MS MARCO Passage(msmarco). This is a set of web search queries from Bing with associated passages. We use the passage-ranking subset and sample 1M training pairs and 2K evaluation queries from the official train/test splits. Unlike NQ, we do not enforce the “seen-document” constraint on MS MARCO (because enforcing it would shrink the evaluation set to only few hundred queries).
Models used for comparison. We implement two representative systems for both DR and GR and intentionally avoid complex variants to keep comparisons fair and transparent. For DR, we implement (i) a standard bi-encoder in the spirit of DPR (dpr) with inner-product scoring; and (ii) a multi-vector late-interaction variant like ColBERT v1 (colbert). For GR, we implement two varying about the docid design and train/inference follow the DSI-style (dsi): (i) codebook docidsbuilt via residual quantization, each docid is a length-6 sequence of 8-bit code indices; and (ii) textual docidsthat directly use the title as the document identifier. All GR decoding is prefix-constrained by a trie constructed from the set of valid docids.
@@
Training and inference. To control for capacity and pretraining, all DR models are built on Qwen3-Embedding-0.6B, and all GR models use Qwen3-0.6B (qwen3). Unless otherwise noted, we train with the Adam optimizer (adam) using its default settings. At inference time, DR retrieves top-$k$ candidates using FAISS-based ANN search (ance), while GR performs top-$k$ constrained decoding over the docid trie.